# Bounding causal effects with an unknown mixture of informative and non-informative missingness

### v3

@@ 3.1 General bounds
Under Assumptions 1-5, the ATE $\Psi_{0}$ is not identifiable due to the presence of informative missingness by $U_{NI}$. We can, however, still bound the ATE using only quantities from the observed data under no additional assumptions. We state this result in Proposition 2.
@@
Proposition 2 follows by the law of iterated expectations and the fact that $Y$ is binary (or bounded and rescaled). This result has also been noted, in, for example, gabriel2023nonparametric; mattei2014identification. These bounds are sharp because they can be attained in the case where $\pi_{a}^{*}(\mathbf{x})=\pi_{a}(\mathbf{x})$ and $\mu^{\star}_{a}(\mathbf{x})=a$ (for the upper bound) or $\mu^{\star}_{a}(\mathbf{x})=1-a$ (for the lower bound). Proofs for this and all subsequent propositions are available in Appendix 9.
Consider a simple example with a single binary covariate $X$ with $\mathbb{P}(X=1)=0.7$, where the outcome risk is twice as high in the informatively missing versus non-informatively missing groups for either treatment arm ($\mu^{*}_{a}(x)=2\mu_{a}(x)$), the treatment doubles the risk of the outcome (so that $\mu_{1}(x)=2\mu_{0}(x)$), and $\mu_{0}(x)=0.1+0.05x$. Assume there is a moderate amount of informative missingness with $\pi_{0}^{\star}(0)=0.07,\pi_{0}^{\star}(1)=0.12,\pi_{1}^{\star}(0)=0.14$, and $\pi_{1}^{\star}(1)=0.19$. For simplicity, assume that all the non-informative missingness has half of these same probabilities ($\pi_{a}^{*}(x)=\frac{2}{3}\pi_{a}(x)$). Treatment is randomly assigned ($e(x)=0.5,x=0,1$). This implies that approximately 20% of the outcomes are missing, with approximately 70% of the missingness informative.
In this scenario, the true ATE is $\Psi_{0}=0.17$. However, the naive estimand – which assumes that all missingness is non-informative – is ${\widetilde{\Psi}}=0.14$, approximately 20% less. Plugging into (LABEL:assumption-free-bounds), we can bound the ATE in the range $\ell_{0},u_{0}]=[-0.06,0.34]$. This bound is somewhat wide, and it leaves open the possibility that the true causal effect is negative, even though the naive estimate is sizeable and positive. We will next outline additional assumptions to construct more informative bounds.
@@ 3.2 Narrower bounds
$\mathbb{P}\left\{\frac{\pi_{a}^{*}(\mathbf{X})}{\pi_{a}(\mathbf{X})}=\delta_{a}\right\}=1\text{ for some }\delta_{a}\in(0,1)$.
@@
$\mathbb{P}\left\{\delta_{a\ell}\leq\frac{\pi_{a}^{*}(\mathbf{X})}{\pi_{a}(\mathbf{X})}\leq\delta_{au}\right\}=1\text{ for some }\delta_{a\ell},\delta_{au}\in[0,1]$.
@@
Under Assumptions 1-5, and either 6.b or 6.a, $\ell_{0,\delta}\leq\Psi_{0}\leq u_{0,\delta}$, where
$\ell_{0,\delta}=\tilde{\Psi}-\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})+\delta_{0u}\pi_{0}(\mathbf{X})\{1-\mu_{0}(\mathbf{X})\}\right]$ (3)
$u_{0,\delta}=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\{1-\mu_{1}(\mathbf{X})\}+\delta_{0u}\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right].$ (4)
Researchers do not in principle ever need to reason about $\delta_{au}$ if they do not want or feel capable to. The bounds in equation (LABEL:assumption-free-bounds) may be recovered by simply setting $\delta_{au}=1$ to obtain the widest possible bounds. However, in many applications, subject-area experts and analysts may be able to reason about the proportion of non-informative missingness by, for example, examining what proportion of missing observations are missing due to the study ending prior to obtaining their outcome measurement, rather than missing for another reason. Alternatively, it may be possible to use auxiliary data to inform possible values of these parameters. For example, with access to an auxiliary dataset where $U_{I}$ is observed, it may be possible to estimate the ratio $\mathbb{P}(U_{I}=1)/\mathbb{P}(C=1)$, $\mathbb{P}(U_{I}=1\mid\mathbf{x})/\mathbb{P}(C=1\mid\mathbf{x})$, or even $\pi_{a}^{\star}(\mathbf{x})$ itself, depending on the available data elements, and to use these to inform choices of $\delta_{au}$. Alternatively, a dataset may contain some covariate $V$ where it is believed that $\mathbb{P}(V=1\mid\mathbf{x})/\mathbb{P}(C=0\mid\mathbf{x})=\beta\pi^{\star}_{a}(\mathbf{x})$ for some constant $\beta$. For example, $U_{I}$ is death, $V$ may be some kind of surgery with a similar ($\beta=1$) prevalence in most population strata defined by covariates $\mathbf{x}$.
In addition, we may also invoke assumptions that relate $\mu_{a}^{\star}$ to $\mu_{a}$ to further narrow these bounds. First, consider the following monotonicity constraints, which result in the bounds given in Proposition 3.
Assumption 7.c (Monotonicity, positive).
$\mathbb{P}\{\mu_{a}^{*}(\mathbf{X})\geq\mu_{a}(\mathbf{X})\}=1,\qquad a=0,1$.
Assumption 7.d (Monotonicity, negative).
$\mathbb{P}\{\mu_{a}^{*}(\mathbf{X})\leq\mu_{a}(\mathbf{X})\}=1,\qquad a=0,1$.
@@
Under Assumptions 1-5, 7.c, and either 6.a or 6.b, $\ell^{\text{pos}}_{0}\leq\Psi_{0}\leq u^{\text{pos}}_{0}$, where
@@
Alternatively, under Assumptions 1-5, 7.d, and 6.b, we may obtain the bounds $\ell^{\text{neg}}_{0}\leq\Psi_{0}\leq u^{\text{neg}}_{0}$, where
$\ell^{\text{neg}}_{0}$ $=\tilde{\Psi}-\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})\}\right],$
@@
Consider again the example in Section 3.1, where we obtained the general bounds [0.06,0.34][-0.06,0.34]. Under positive monotonicity alone, which holds in our example, we can obtain the slightly more narrow bounds $[0.01,0.31]$ (via Proposition 3). If we are further willing to assume that the proportion of informative missingness in either treatment arm is at most 80% ($\delta_{1u}=\delta_{0u}=0.8$), we can further refine the bounds $[0.03,0.28]$.
Typically, the informative missingness that most threatens identification is monotonic – e.g., sicker patients are less likely to have complete data – but we do not assume that monotonicity assumptions will be realistic for all applications. For example, patients may be less likely to follow up with healthcare providers if they are feeling healthier (and thus not in immediate need of care) as well as if they are too sick to seek care. As an alternative, we may consider assumption 7.e below. [^5] Implicit in assumption 7.e and 7.f is that $\mu_{a}(\mathbf{x})>0$ for all $a,\mathbf{x}$. Alternatively, one could re-state this assumption on the support of $\mathbf{x}$ where $\mu_{a}(\mathbf{x})>0$, and define $\mu_{a}^{\star}(\mathbf{x})=0$ where $\mu_{a}(\mathbf{x})=0$. It would then be straightforward as below to derive analogous expressions and estimators for the implied bounds, though we do not explore this further in this paper.
Assumption 7.e (Bounded outcome risk).
@@
This assumptions restricts how much informative missingness may increase/decrease the risk of the primary outcome, resulting in the following bound:
@@
Under Assumptions 1-5, 7.e, and 6.b, $\widetilde{\ell}\leq\Psi_{0}\leq\widetilde{u}$, where
$\widetilde{\ell}$ $=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}(\tau_{1}^{-1}-1)\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0u}(\tau_{0}-1)\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right],$
$\widetilde{u}$ $=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}(\tau_{1}-1)\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0u}(\tau_{0}^{-1}-1)\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right].$
@@
This bound is sharp when $\tau_{a}\mu_{a}(\mathbf{x})\leq 1$ for $a=0,1$. However, this may not hold for any given choice of $\tau_{a}$. This will be a greater concern when the risk of the outcome is high (near one), leading to an overly conservative bound. One may instead start from the assumption that $\frac{\mu_{a}^{\star}(\mathbf{x})}{\mu_{a}(\mathbf{x})}\leq\min(1/\mu_{a}(\mathbf{x}),\tau_{a})$ and integrate over the resulting expression to obtain a sharp bound. We discuss this further in Remark 10 in Appendix 9.2.
@@
Continuing with our example, assuming that the risk of the outcome in observations with missing outcomes is no more than three times and no less than one-third of the risk of the outcome in the non-censored observations ($\tau=3$), then even if we allow all missingness to be informative ($\delta_{au}=1$), we obtain the bounds $[0.04,0.28]$. If we further restrict the proportion of informative missingness to be no more than 80%, we obtain the slightly narrower bounds $[0.06,0.25]$.
We may again use either subject-area expertise or a data-driven approach to determine reasonable values of $\tau_{a}$ for a given application. For example, researchers may know that the prevalence of some outcome is not more than twice as likely under informative missingness, or less than half as likely in either treatment arm (i.e. $\tau_{a}=2$ for $a=0,1$). As a possible data driven approach, one may assume a model where there exist covariates $(X,U_{X})$ such that $Y\perp U_{I}\mid X,U_{X}$. This model precludes the possibility that $U_{I}$ affects $Y$ directly; however, it does allow us to calibrate the confounding effects of $U_{I}$ with respect to the observed covariates. To be precise, we can compute the following quantity
$\Omega_{a}$ $=\frac{\mathbb{E}[Y\mid A=a,C=1]}{\mathbb{E}[Y\mid A=a,C=0]}$
$=\frac{\mathbb{E}[\mathbb{E}[Y\mid A=a,C=0,\mathbf{x}]\mid A=a,C=1]}{\mathbb{E}[\mathbb{E}[Y\mid A=a,C=0]\mid A=a,C=1]}$
which provides an estimate of the ability of the entire covariate set to control for informative missingness within treatment arm $A=a$ under the assumption that all missingness is non-informative conditional on the covariates – $Y\perp C\mid X,A$. This may give researchers a heuristic for calibrating what types of outcome risk increases/decreases are reasonable. That said, this is simply a starting approach. Determining methods to better calibrate this sensitivity parameter with respect to specific variables rather than the entire set, or extending the calibration method to include cases where $U_{I}$ directly effects $Y$, would be an interesting avenue for future research.
3.3 Point identification and sensitivity analysis
@@
The preceding sections have provided bounds for the ATE as a function (in many cases) of parameters encoding the amount and nature of informative missingness. These bounds are most useful in settings where these parameters may be specified, ascertained, or guessed. In other situations, one may prefer to start from the observed naive association $\widetilde{\Psi}$} and instead ask what type of informative missingness would explain it away. Each of the previously proposed bounds may be “inverted” to obtain such a sensitivity analysis. We may return to the running example to see this. In place of the bounds based on Proposition 3 – previously given in Section 3.2 as [0.03, 0.28] with $\delta_{au}=0.8$ – we may rearrange terms in Proposition 3 to determine what value of $\delta_{0u}$ or $\delta_{1u}$ are required for $\ell_{0}^{\text{pos}}>0$ or $u_{0}^{\text{pos}}<0$. In this example, we can guarantee that $\ell_{0}^{\text{pos}}\geq 0$ if $\delta_{1u}\leq{\widetilde{\Psi}}/\mathbb{E}\left[\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})\}\right]=1.08$ (which is always true because $\delta_{1u}\leq 1$ by definition). This type of analysis may be appealing in settings where researchers struggle to specify particular values of the sensitivity parameters but yet have a sense of which sensitivity parameters would be implausible.
@@
Furthermore, in order to ease communication of such analyses to collaborators and non-statisticians, one may wish to reduce the complexity of such an analysis. In this section, we propose such a simplified analysis based primarily on a single sensitivity parameter $\tau$ that indicates how much informative missingness increases risk.
To operationalize this simplified analysis, we first show how $\Psi_{0}$ may be identified by assuming that the proportion of informative missingness ratio and outcome risk ratio are constant and known quantities. Assumption 7.f and Proposition 5 formalize this result.
Assumption 7.f (Known outcome risk ratio).
@@
Proposition 5 (Point identification of ATE).
Under Assumptions 1-5, 7.f, and 6.a,
$\Psi_{0}=\tilde{\Psi}+\mathbb{E}\left[(\tau-1)\left\{\delta_{1}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0}\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right\}]\right].$
@@
We can use Proposition 5 to examine what combination of $(\tau,\delta_{0},\delta_{1})$ would be required to explain away the observed naive association $\widetilde{\Psi}$}, we may make quite succinct statements based only on $\tau$ when we let $\delta_{0},\delta_{1}$ take extreme values. For example, assume, as in our running example, that ${\widetilde{\Psi}}>0$. Then $\tau$ must be at least as large as $1+{\widetilde{\Psi}}/\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]$ to possibly explain away the observed association – that is, if $\tau<1+{\widetilde{\Psi}}/\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]$, then $\Psi_{0}>0$ under the assumptions of Proposition 5. In the example, we have ${\widetilde{\Psi}}=0.14,\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]=0.02,\mathbb{E}[\mu_{1}(\mathbf{X})\pi_{1}(\mathbf{X})]=0.06$. Therefore, unless informative missingness increases risk by at least a factor of $\tau=7.62$, there is no way to explain the observed naive association of ${\widetilde{\Psi}}=0.14$. And, even if $\tau$ could reasonably be as large as that, the proportion of informative missingness in the control group $\delta_{0}$ has to be at least as large as $\frac{\delta_{1}\mathbb{E}[\mu_{1}(\mathbf{X})\pi_{1}(\mathbf{X})]+{\widetilde{\Psi}}/(\tau-1)}{\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]}$ when $\tau\neq 1$. If we take $\tau=10$ in the example, then we require $\delta_{0}>0.74+3.1\delta_{1}$, which restricts the amount of informative missingness in the treated group to be essentially nonexistent and in the control group to be nearly universal. We plot the regions implied by Proposition 5 in Figure 1. For a given value of $\tau$, corresponding to one line in Figure 1, $(\delta_{0},\delta_{1})$ must lie in the region to the lower right of the line if informative missingness could possibly explain the observed association.
@@
Remark 4.
The sensitivity analysis in this section is constructed to be as simple as possible. To this end, this sensitivity analysis relies on Assumption 7.f which states that the increased risk due to informative missingness is the same in the treated and control groups. This is not required, and one is free to instead assume two different risk ratios $\tau_{0},\tau_{1}$. This additional flexibility comes at the expense of a sensitivity analysis that is harder to interpret and therefore prefer the former approach. Furthermore, all of the bounds in preceding sections may be “inverted” to make similar statements about what type of informative missingness might explain away the observed association under their stated assumptions. We think the analysis in this section is the easiest to communicate.

### v4

@@ 3.1 General bounds
Under Assumptions 1-5, the ATE $\Psi_{0}$ is not identifiable due to the presence of informative missingness by $U_{I}$. We can, however, still bound the ATE using only quantities from the observed data under no additional assumptions. We state this result in Proposition 2.
@@
Proposition 2 follows by the law of iterated expectations and the fact that $Y$ is binary (or bounded and rescaled). This result has also been noted, in, for example, Mattei et al. (2014); Gabriel et al. (2023). These bounds are sharp because they can be attained in the case where $\pi_{a}^{\star}(\mathbf{x})=\pi_{a}(\mathbf{x})$ and $\mu^{\star}_{a}(\mathbf{x})=a$ for all $\mathbf{x}$ (for the upper bound) or $\mu^{\star}_{a}(\mathbf{x})=1-a$ for all $\mathbf{x}$ (for the lower bound). Proofs for this and all subsequent propositions are available in Appendix 10.
Consider a simple example with a single binary covariate $X$ with $\mathbb{P}(X=1)=0.7$, where the outcome risk is twice as high in the informatively missing versus non-informatively missing groups for either treatment arm ($\mu^{*}_{a}(x)=2\mu_{a}(x)$), the treatment doubles the risk of the outcome (so that $\mu_{1}(x)=2\mu_{0}(x)$), and $\mu_{0}(x)=0.1+0.05x$. Assume there is a moderate amount of informative missingness with $\pi_{0}^{\star}(0)=0.07,\pi_{0}^{\star}(1)=0.12,\pi_{1}^{\star}(0)=0.14$, and $\pi_{1}^{\star}(1)=0.19$. For simplicity, assume that all the non-informative missingness has half of these same probabilities (implying that $\pi_{a}^{\star}(x)=\frac{2}{3}\pi_{a}(x)$). Treatment is randomly assigned ($e(x)=0.5,x=0,1$). This implies that approximately 21% of the outcomes are missing, with two-thirds of the missingness informative.
In this scenario, the true ATE is $\Psi_{0}\approx 0.17$. However, the naive estimand – which assumes that all missingness is non-informative – is ${\widetilde{\Psi}}\approx 0.135$, approximately 20% less. Plugging into (LABEL:assumption-free-bounds), we can bound the ATE in the range $\ell_{0},u_{0}]\approx[-0.07,0.35]$. This bound is somewhat wide, and it leaves open the possibility that the true causal effect is negative, even though the naive estimate is sizeable and positive. We will next outline additional assumptions to construct narrower bounds.
@@ 3.2 Narrower bounds
$\mathbb{P}\left\{\frac{\pi_{a}^{\star}(\mathbf{X})}{\pi_{a}(\mathbf{X})}=\delta_{a}\right\}=1\text{ for some }\delta_{a}\in[0,1]$.
@@
$\mathbb{P}\left\{\delta_{a\ell}\leq\frac{\pi_{a}^{\star}(\mathbf{X})}{\pi_{a}(\mathbf{X})}\leq\delta_{au}\right\}=1\text{ for some }\delta_{a\ell},\delta_{au}\in[0,1]$.
@@
Under Assumptions 1-5, and either 6.a or 6.b, $\ell_{0}^{\star}\leq\Psi_{0}\leq u_{0}^{\star}$, where
$\ell_{0}^{\star}=\tilde{\Psi}-\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})+\delta_{0u}\pi_{0}(\mathbf{X})\{1-\mu_{0}(\mathbf{X})\}\right]$ (3)
$u_{0}^{\star}=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\{1-\mu_{1}(\mathbf{X})\}+\delta_{0u}\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right].$ (4)
We may also invoke assumptions that relate $\mu_{a}^{\star}$ to $\mu_{a}$ to further narrow these bounds. First, consider the following monotonicity constraints, which result in the bounds given in Proposition 3.
Assumption 7.a (Monotonicity, positive).
$\mathbb{P}\{\mu_{a}^{\star}(\mathbf{X})\geq\mu_{a}(\mathbf{X})\}=1,\qquad a=0,1$.
Assumption 7.b (Monotonicity, negative).
$\mathbb{P}\{\mu_{a}^{\star}(\mathbf{X})\leq\mu_{a}(\mathbf{X})\}=1,\qquad a=0,1$.
@@
Under Assumptions 1-5, 7.a, and either 6.a or 6.b, $\ell^{\text{pos}}_{0}\leq\Psi_{0}\leq u^{\text{pos}}_{0}$, where
@@
Alternatively, under Assumptions 1-5, 7.b, either 6.a or 6.b, we may obtain the bounds $\ell^{\text{neg}}_{0}\leq\Psi_{0}\leq u^{\text{neg}}_{0}$, where
$\ell^{\text{neg}}_{0}$ $=\tilde{\Psi}-\mathbb{E}\left[\delta_{1u}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})\right],$
@@
Consider again the example in Section 3.1, where we obtained the general bounds [0.07,0.35][-0.07,0.35]. Under positive monotonicity alone, which holds in our example, we can obtain the slightly more narrow bounds [0.00,0.33][-0.00,0.33] (via Proposition 3). If we are further willing to assume that the proportion of informative missingness in either treatment arm is at most 80% ($\delta_{1u}=\delta_{0u}=0.8$), we can further refine the bounds to be approximately $[0.03,0.29]$.
Typically, the informative missingness that most threatens identification is monotonic – e.g., sicker patients are less likely to have complete data – but we do not assume that monotonicity assumptions will be realistic for all applications. For example, patients may be less likely to follow up with healthcare providers if they are feeling healthier (and thus not in immediate need of care) as well as if they are too sick to seek care. As an alternative, we may consider assumption 7.c below. [^5] Implicit in assumption 7.c and 7.d is that $\mu_{a}(\mathbf{x})>0$ for all $a,\mathbf{x}$. Alternatively, one could re-state this assumption on the support of $\mathbf{x}$ where $\mu_{a}(\mathbf{x})>0$, and define $\mu_{a}^{\star}(\mathbf{x})=0$ where $\mu_{a}(\mathbf{x})=0$. It would then be straightforward as below to derive analogous expressions and estimators for the implied bounds, though we do not explore this further in this paper.
Assumption 7.c (Bounded outcome risk).
@@
This assumption restricts how much informative missingness may increase/decrease the risk of the primary outcome, resulting in the following bound:
@@
Under Assumptions 1-5, either 6.b or 6.a, and 7.c, $\widetilde{\ell}_{0}\leq\Psi_{0}\leq\widetilde{u}_{0}$, where
$\widetilde{\ell}_{0}$ $=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}(\tau_{1}^{-1}-1)\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0u}(\tau_{0}-1)\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right],$
$\widetilde{u}_{0}$ $=\tilde{\Psi}+\mathbb{E}\left[\delta_{1u}(\tau_{1}-1)\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0u}(\tau_{0}^{-1}-1)\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right].$
@@
This bound is sharp when $\tau_{a}\mu_{a}(\mathbf{x})\leq 1$ for $a=0,1$ and all $\mathbf{x}$. However, this may not hold for any given choice of $\tau_{a}$. This will be a greater concern when the risk of the outcome is high (near one), leading to an overly conservative bound. One may instead start from the assumption that $\frac{\mu_{a}^{\star}(\mathbf{x})}{\mu_{a}(\mathbf{x})}\leq\min(1/\mu_{a}(\mathbf{x}),\tau_{a})$ and integrate over the resulting expression to obtain a sharp bound. We give the explicit formulation of the resulting bounds, along with doubly-robust style estimators of these quantities, in Appendix 9.
@@
Continuing with our example, assuming that the risk of the outcome in observations with missing outcomes is no more than three times and no less than one-third of the risk of the outcome in the non-censored observations ($\tau=3$), then even if we allow all missingness to be informative ($\delta_{au}=1$), we obtain the bounds $[0.04,0.29]$. If we further restrict the proportion of informative missingness to be no more than 80%, we obtain the slightly narrower bounds $[0.06,0.26]$.
3.3 Choosing sensitivity parameters
Some of our proposed sensitivity analyses require specifying values for unknown parameters, such as the proportion of informative missingness $\delta_{au}$ and the outcome risk ratio $\tau_{a}$. Depending on the context of the application, we recommend four broad strategies to choose these values: (1) imposing the weakest assumptions possible; (2) using subject-area expertise; (3) using data-informed approaches; and (4) conducting an analysis across a range or grid of parameters (which, to be clear, is not mutually exclusive with the first three strategies). We discuss these strategies in more detail below.
First, we consider minimal assumption analyses. Researchers who feel unable to reason about the values of these parameters may prefer to make as few assumptions as possible. In this case, one may set $\delta_{au}=1$ (and, where relevant, $\delta_{a\ell}=0$). This corresponds to allowing the maximum possible proportion of informative missingness and therefore imposes no restriction on the missingness mechanism. Bounds obtained under these values represent the most conservative conclusions that can be drawn under the assumed structural conditions. Similarly, if we wish to make no assumptions on $\tau_{a}$, researchers may use the general bounds provided in Proposition 2. Naturally, these bounds will typically be the widest.
Second, domain knowledge can meaningfully constrain the sensitivity parameters for some applications. For example, investigators may know that some missing outcomes are due to administrative censoring (e.g., the study ending before outcomes are observed), which would suggest that only a fraction of missingness is plausibly informative. Such reasoning can justify values of $\delta_{au}<1$. Similarly, subject-matter experts may have intuition about plausible outcome risk differences under informative missingness. For example, investigators might believe that the outcome risk among informatively missing individuals could plausibly be no more than twice as large (or half as large) as among observed individuals, corresponding to $\tau_{a}=2$.
Third, we may use data to help calibrate the sensitivity parameters. For example, in some cases an auxiliary dataset may be available that contains information on the informative missingness indicator $U_{I}$, allowing one to estimate quantities such as $\frac{\mathbb{P}(U_{I}=1)}{\mathbb{P}(C=1)}$, or even $\pi_{a}^{\star}(\mathbf{X})$ directly, depending on what variables are observed. These quantities can provide empirical guidance about plausible ranges for $\delta_{au}$. To inform $\tau_{a}$, even with no auxiliary dataset, one may compare the predictive strength of observed covariates with respect to missingness. Suppose there exist covariates $(\mathbf{X},U_{X})$ such that
$Y\perp U_{I}\mid\mathbf{X},U_{X}.$
Although this assumption rules out direct causal effects of $U_{I}$ on $Y$, it allows the observed covariates to serve as a proxy for the confounding strength of informative missingness. One can then compute
$\Gamma_{a}=\frac{\mathbb{E}[Y\mid A=a,C=1]}{\mathbb{E}[Y\mid A=a,C=0]}=\frac{\mathbb{E}[\mathbb{E}[Y\mid A=a,C=0,\mathbf{X}]\mid A=a,C=1]}{\mathbb{E}[\mathbb{E}[Y\mid A=a,C=0,\mathbf{X}]\mid A=a,C=0]},$
which summarizes the ability of the entire observed covariate set to explain outcome differences between censored and uncensored individuals within treatment arm $A=a$. This quantity can serve as a heuristic benchmark when specifying values of $\tau_{a}$. While these proposals outline some basic ideas, developing principled methods for calibrating these sensitivity parameters to observed data would be a useful area for future work.
Finally, regardless of whether researchers wish to make minimal assumptions, use subject-area expertise, or a data-driven approach, they may also wish to evaluate the bounds over a range or a grid of sensitivity parameter values. For example, subject-area expertise may only inform the range of plausible $\delta_{au}$; similarly, a calibration analysis may suggest a value for $\mathbb{P}(U_{I}=1)/\mathbb{P}(C=1)$, which can restrict the plausible range of $\pi_{a}^{\star}(\mathbf{x})$, but does not inform $\delta_{au}$ directly. In these cases one may evaluate the bounds across a range or grid, and plot how the bounds vary over $(\delta_{au},\tau_{a})$ to provide a transparent view of how conclusions depend on the assumed parameters. This approach is particularly useful when analysts cannot commit to a single parameter choice but wish to understand how robust their conclusions are across a range of plausible assumptions.
3.4 Point identification and tipping point sensitivity analyses
The strategies outlined above all require specifying values or ranges for the sensitivity parameters. An alternative approach is to instead ask what magnitude of informative missingness would be required to explain away the observed association. This type of analysis, often called a “tipping-point” analysis, starts from the naive estimate $\tilde{\Psi}$ and determines the values of the sensitivity parameters for which the causal effect could change sign. Each of the previously proposed bounds may be “inverted” to obtain such a sensitivity analysis. That is, for fixed values of all but one sensitivity parameter, one may set the expression equal to zero and solve for the value of the parameter that would satisfy the equation. However, we emphasize that the grid searches proposed above also often reveal tipping points without directly solving for these values, and should not be viewed as incompatible or mutually exclusive with these tipping point analyses.
@@ 3.4 Point identification and tipping point sensitivity analyses
Nevertheless, consider again the running example above to illustrate. In place of the bounds based on Proposition 3, we may rearrange terms in Proposition 3 to determine what value of $\delta_{0u}$ or $\delta_{1u}$ are required for $\ell_{0}^{\text{pos}}>0$ or $u_{0}^{\text{pos}}<0$. In this example, we can guarantee that $\ell_{0}^{\text{pos}}\geq 0$ if $\delta_{0u}\leq{\widetilde{\Psi}}/\mathbb{E}\left[\pi_{0}(\mathbf{X})(1-\mu_{0}(\mathbf{X}))\right]\approx 1.00$ (which is always true because $\delta_{0u}\leq 1$ by definition). This type of analysis may be appealing in settings where researchers struggle to specify particular values of the sensitivity parameters but yet have a sense of which sensitivity parameters would be implausible.
@@
However, in cases where we also wish to make more substantive restrictions on the outcome risk ratio, the sensitivity analysis proposed in Proposition 4 may be harder to invert in an interpretable way, in part due to the numbers of parameters. In order to ease communication of such analyses to collaborators and non-statisticians, one may wish to reduce the complexity of this analysis. We therefore also propose a simplified analysis based primarily on a single sensitivity parameter $\tau$ that indicates how much informative missingness changes the risk. Specifically, we first show how $\Psi_{0}$ may be identified by assuming that the proportion of informative missingness ratio and outcome risk ratio are constant and known quantities. Assumption 7.d and Proposition 5 formalize this result.
Assumption 7.d (Known outcome risk ratio).
@@
Proposition 5 (Point identification of the ATE).
Under Assumptions 1-5, 6.a, and 7.d,
$\Psi_{0}=\tilde{\Psi}+\mathbb{E}\left[(\tau-1)\left\{\delta_{1}\pi_{1}(\mathbf{X})\mu_{1}(\mathbf{X})-\delta_{0}\pi_{0}(\mathbf{X})\mu_{0}(\mathbf{X})\right\}\right].$
@@
We can then use Proposition 5 to examine what combination of $(\tau,\delta_{0},\delta_{1})$ would be required to explain away the observed naive association $\widetilde{\Psi}$}, making succinct statements based only on $\tau$ when we let $\delta_{0},\delta_{1}$ take extreme values. For example, assume, as in our running example, that ${\widetilde{\Psi}}>0$. Then, setting $\delta_{0}=1$ and $\delta_{1}=0$ (the most adversarial configuration of informative missingness), we may conclude that $\tau$ must be at least as large as $1+{\widetilde{\Psi}}/\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]$ to possibly explain away the observed association. Correspondingly, if $\tau<1+{\widetilde{\Psi}}/\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]$, then $\Psi_{0}>0$ under the assumptions of Proposition 5.
In our running example, this analysis allows us to conclude that unless informative missingness increases risk by at least a factor of $\tau=7.12$, there is no way to explain the observed naive association of ${\widetilde{\Psi}}=0.135$. And, even if $\tau$ could reasonably be as large as that, the proportion of informative missingness in the control group $\delta_{0}$ has to be at least as large as $\frac{\delta_{1}\mathbb{E}[\mu_{1}(\mathbf{X})\pi_{1}(\mathbf{X})]+{\widetilde{\Psi}}/(\tau-1)}{\mathbb{E}[\mu_{0}(\mathbf{X})\pi_{0}(\mathbf{X})]}$ when $\tau\neq 1$. If we take $\tau=10$, then we require $\delta_{0}>0.68+3.29\delta_{1}$, which restricts the amount of informative missingness in the treated group to be very small and in the control group to be nearly universal. We plot the regions implied by Proposition 5 in Figure 1. For a given value of $\tau$, corresponding to one line in Figure 1, $(\delta_{0},\delta_{1})$ must lie in the region to the lower right of the line if informative missingness could possibly explain the observed association.
@@
We again construct this sensitivity analysis to be as simple as possible. To this end, this sensitivity analysis relies on Assumption 7.d which states that the increased risk due to informative missingness is the same in the treated and control groups. This is not required, and one may instead assume two different risk ratios $\tau_{0},\tau_{1}$. This additional flexibility comes at the expense of a sensitivity analysis that is harder to interpret, and we therefore prefer the former approach. Furthermore, we again stress that all of the bounds in preceding sections may be “inverted” to make similar statements about what type of informative missingness might explain away the observed association under their stated assumptions. However, the analysis proposed in this section may be relatively easier to communicate than others.